# See2Refine: Vision-Language Feedback Improves LLM-Based eHMI Action Designers

### v1

@@ 3.5.1 Importance-based Scenario Sampling
For each scenario $i$, the shared action database stores tuples $(i,m,K_{i}^{m})$, where $m$ indexes the source model/round and $K_{i}^{m}$ is the kernel score. From these records, we compute $K_{i}^{\text{best}}=\max_{m}K_{i}^{m}$, $K_{i}^{\text{worst}}=\min_{m}K_{i}^{m}$, $\Delta K_{i}=K_{i}^{\text{best}}-K_{i}^{\text{worst}}$, and the candidate count $N_{i}$. We define the importance score:
@@
  • Low best scores $K_{i}^{\text{best}}$ (room for improvement)
  • Small gaps $\Delta K_{i}$ (limited diversity)
  • High worst scores $K_{i}^{\text{worst}}$ (easier to improve)
  • Fewer previous sampling attempts
@@ 3.5.2 Diverse Action Generation
For each sampled scenario, we generate diverse eHMI action sequences with the current fine-tuned model using diverse beam search Vijayakumar et al. (2016) (six beam groups with one beam each). Each output is validated against modality-specific JSON schemas (eyes/arm/lightbar) via Pydantic, and invalid formats are discarded. We then render and evaluate the valid actions using the same pipeline as the initial database, compute their kernel scores, and append them to the shared action database. Repeating this step iteratively expands the candidate set and improves both diversity and quality over training rounds.
@@ 3.5.3 Preference Pair Construction
After expanding the shared score database, we form DPO preference pairs $(y^{+},y^{-})$ per scenario, where $y^{+}$ is a higher-quality action than $y^{-}$. This setup is analogous to human preference ordering, as discussed in prior work Xia et al. (2025). Pair extraction is two-stage. (1) Max–min: for each scenario $i$, select the highest and lowest-scoring actions,
@@ 4.1 Experimental Settings
All finetuning processes use ShareGPT-formatted data Chiang et al. (2023) with an 80/20 train/test split. We fine-tune with LoRA Hu et al. (2022) in bfloat16: 3 epochs for Format-aware Fine-tuning and 1 epoch for DPO (to reduce overfitting, similar to prior works Bai et al. (2022b); Touvron et al. (2023); Liu et al. (2024)). DPO uses weight=1, sigmoid loss, and a preference coefficient of 0.01.
@@
VLM rating achieves approximately 90 evaluations per second with 8 threads. Single-thread inference requires about 7 seconds per scenario for the 7B model and 6 seconds per scenario for the 1.5B model. The rendering takes roughly 45 seconds per clip, operating in batch parallelization. Fine-tuning duration is approximately 3 hours for the 7B model and 1 hour 10 minutes for the 1.5B model in Format-aware Fine-tuning, and around 5 hours (7B) or 3 hours 30 minutes (1.5B) per DPO round. Excluding dataset construction and assuming three DPO rounds, the total co-learning time amounts to at least 60 hours for the 7B model and 38 hours for the 1.5B model.
@@ 4.2.3 Metric-wise Analysis
Metric choices are inherited from prior eHMI work, but their effectiveness for VLM raters is unclear. Therefore, we analyze the VLM score distributions for three milestones: (1) the initial action database, (2) DesignerLLM-Apprentice-7B, and (3) DesignerLLM-7B. As illustrated in Figure 7 in the Appendix, metrics behave differently over training. Phase 2 metrics (User Acceptance, Consistency) improve most clearly: their means increase and distributions tighten, suggesting that they are the most sensitive and reliable signals for VLM-based evaluation. In contrast, Phase 1 metrics show weaker trends; targeting improves slightly, while trust changes little and provides poor preference separation in the initial database, making it less useful for forming training pairs. This likely reflects a mismatch between VLM- and human-perceived trust in our setting.
@@ 4.3 Human Preference Alignment
We conducted a user study to (1) test whether improvements transfer from VLM raters to people, and (2) examine potential gaps between VLM and human preferences. Details are in Appendix A.
Setup. We evaluated two common modalities (eyes, lightbar) on eight scenario-message pairs from prior work Xia et al. (2025). We compared five models: DesignerLLM-base, DesignerLLM, and three commercial LLMs (OpenAI GPT-5, Claude Opus 4.1, Grok 4). Each model generated actions that were rendered as 1080p videos at 12 FPS with full scene context (80 videos total).
Participants and measures. We recruited 18 participants (21–30 years, M=26M{=}26, SD=2.7SD{=}2.7), split evenly by modality. For each scenario, participants rated videos in two stages: without the intended message (targeting, trust, perceived safety, mental workload) and with the intended message revealed (acceptance, consistency).
Results. Across both modalities, DesignerLLM consistently outperformed the base model and ranked 4th overall (Figure 4). Relative to the base model, DesignerLLM improved average scores by 7.7% (eyes) and 7.9% (lightbar), with the largest gains in targeting (15.4% eyes; 12.6% lightbar) and acceptance (5.0% eyes; 12.5% lightbar). Commercial models achieved higher absolute ratings, but the improvements from our framework were statistically meaningful for most metrics, indicating that VLM-aligned training also yields measurable benefits for human-perceived eHMI quality.
@@ 5 Conclusion
We introduce See2Refine leveraging VLM perceptual feedback as an automated approach to progressively enhance the capabilities of LLM-based eHMI action designers. We perform extensive experiments and explore methods for effectively expanding the action dataset while preserving alignment with VLM preference performance. Additionally, we conduct a user study demonstrating that insights from VLM raters can also improve human perception scores. Finally, our work serves as an important practice in improving the message delivery capabilities of LLM-actuated systems.
@@
The current setup demonstrates that, with the See2Refine system, LLM-based eHMI action designers can achieve higher human perception scores, despite using VLM as an efficient substitute for real human annotators. However, as noted in previous work Xia et al. (2025), the preference order between a human and VLM is consistent only when the rating score difference between two actions is sufficiently large. Replicating the nuanced preferences of individuals from different cultural backgrounds or age groups remains challenging. Therefore, incorporating a real human annotator into the refinement process would be valuable.
There are two aspects worth exploring. First, effective fine-tuning of VLM rater models could further improve system performance. However, collecting enough data for a single round of fine-tuning is still costly, so the potential marginal gains from such fine-tuning merit careful consideration. Second, rating alone is often not an information-rich method for annotation. If we consider our system as a simulation of a real teacher, employing a teaching approach, using rich cues such as voice, gestures, posture, and other non-verbal signals, could serve as a more natural and effective way to teach LLM designers accurate scenario-action pairs. This approach could also help reduce the ambiguity associated with preference pairs, thereby enabling more effective distillation of human preferences into the LLM training process.
@@
In the current implementation (Section 3.1), we utilize generated scenarios and rendered action videos as visual cues for VLM raters. This approach has proven to be both effective and efficient. However, it primarily focuses on simulating the main elements of typical traffic scenarios, such as automated vehicles, while overlooking other factors that could influence the visual perception outputs for both VLM raters and human road users, such as lighting conditions, obstructions, and other environmental factors. Therefore, the challenge is to incorporate more visual elements that accurately simulate real-world scenarios in rendered videos, while maintaining reasonable rendering times to enhance the robustness of LLM-based action designers in practical applications. One potential trade-off solution is to combine different levels of visual detail within the same training session. For example, including videos with rich visual elements for only a small portion of the training data could improve robustness while keeping overall training costs relatively low.
@@ A.1 Study Design and Procedure
We selected two eHMI modalities (lightbar and eyes) as simpler and common modalities for the user study. Eight traffic scenario–message pairs were adopted from prior work Xia et al. (2025) covering common interactions between AVs and other road users. Five LLMs were tested: two 7B-parameter models (DesignerLLM-base, trained after Format-aware Fine-tuning, and DesignerLLM, after Iterative Preference-based Co-learning) and three state-of-the-art models (OpenAI GPT-5, Claude Opus 4.1, and Grok 4). Each model generated eHMI actions that were rendered into 1080p (1980×1080) videos at 12 FPS to ensure smooth playback for participants in Blender. Unlike the videos provided to VLM raters (Section 3.2.2), these videos include rich contextual information. A total of 80 videos were created. Rendering all scenarios with detailed meshes and visual effects takes approximately 10 hours of computation.
@@
The user study was implemented and deployed using Gradio. Separate interfaces were created for the two eHMI modalities. Each interface consisted of five components: (1) a welcome page, (2) an introduction to general eHMI concepts and the specific modality, including two demonstration videos, (3) a demographic information section, (4) main rating pages for eight scenario–message pairs, and (5) an ending page. The interface included a session-based resume function, allowing participants to continue later using their assigned ID.
@@
In the demographic section, participants anonymously reported their age, gender, and familiarity with eHMIs. For the main user rating pages, eight scenarios are evaluated across different pages, each presented in two stages: the first stage, where the intended messages are not provided, and the second stage, where the intended messages are included. This results in a total of 16 pages. On each page, five videos are shown to participants in a randomized order. As illustrated in Figure 5, each video assessment follows a horizontal layout, with the video on the left and the questionnaires on the right. To ensure no scores are missed, a verification mechanism is in place: participants must complete the current page before proceeding to the next, ensuring the completeness of the questionnaires. Each session took approximately 5 minutes to complete.
@@ A.2 Participants
We recruited 18 participants aged 21 to 30 (mean age: 26, SD = 2.7), consisting of nine males and nine females. They were evenly assigned to evaluate the lightbar and eye modalities. Among the participants, one was highly familiar with eHMIs, possessing extensive knowledge or experience; two had some familiarity, having heard about eHMIs before; and the remaining fifteen were unfamiliar with eHMIs. All participants received a \$10 gift card as compensation.
@@ A.3 Measurements
Without Intended Message. Participants rated their initial impressions of the eHMI using four metrics:
  • Message Targeting (1–9): The extent to which the eHMI message was directed toward the participant.
  • Trust (1–9): The participant’s confidence in and willingness to rely on the eHMI communication.
@@
  • Mental Workload (1–20): The cognitive effort required to interpret the eHMI actions.
With Intended Message. After revealing the intended messages, participants reassessed using two metrics:
  • User Acceptance (1–9): The willingness to accept the eHMI in daily life.
  • Consistency (1–9): The perceived alignment between the displayed eHMI motion and its intended message.
@@ A.4 Result Analysis
Prior to analysis, normality of these participant-level means was assessed with the Shapiro–Wilk test and showed no serious deviations ($p>0.05$) for both modalities. Because the study used a within-subjects design with five model conditions, each metric was analyzed using repeated-measures ANOVA.
@@ A.4.1 Reliability Analysis
We assessed the reliability of the subjective ratings before analyzing model effects. Cronbach’s alpha was used to evaluate internal consistency within each metric across the five models. Internal consistency was excellent for both modalities (eye: $\alpha=0.925$; lightbar: $\alpha=0.953$), with all metrics exceeding $\alpha=0.90$. Inter-rater reliability, estimated with the intraclass correlation coefficient (ICC), was low (eye: $ICC=0.131$; lightbar: $ICC=0.087$), indicating limited absolute agreement across participants. Because our analyses focus on within-subject, relative comparisons in a single session, high internal consistency is the primary requirement; the low ICC limits generalization to settings requiring stable scores across raters or sessions Koo and Li (2016).
@@ A.4.2 Eye Modality Performance
For the eye modality condition with 9 participants, repeated measures ANOVA revealed significant main effects of model type across all six metrics:
Message Targeting. The ANOVA showed a significant main effect of model type ($F(4,32)=5.41$, $p=0.002$, $\eta^{2}=0.403$). Grok 4 ($M=5.47$, $SD=0.89$) and OpenAI GPT-5 ($M=5.47$, $SD=0.88$) achieved the highest ratings, while the base model Qwen2.5-7B-Instruct showed the lowest performance ($M=4.21$, $SD=1.65$). Notably, DesignerLLM ($M=4.86$, $SD=1.21$) demonstrated a 15.4% improvement over the base model, indicating that our co-learning framework successfully enhanced message targeting metric despite its modest 7B parameter size.
Trust. A significant main effect was observed ($F(4,32)=4.66$, $p=0.005$, $\eta^{2}=0.368$). Grok 4 ($M=5.43$, $SD=1.04$) achieved the highest trust ratings, followed by OpenAI GPT-5 ($M=5.32$, $SD=1.15$). DesignerLLM ($M=4.65$, $SD=1.32$) achieved a 7.9% improvement over the base model ($M=4.31$, $SD=1.64$), demonstrating the framework’s effectiveness in generating more trustworthy eHHI action design.
@@
User Acceptance. The ANOVA indicated a significant main effect ($F(4,32)=4.23$, $p=0.007$, $\eta^{2}=0.346$). Grok 4 ($M=5.39$, $SD=1.19$) and OpenAI GPT-5 ($M=5.31$, $SD=1.28$) achieved the highest acceptance ratings. DesignerLLM ($M=4.38$, $SD=1.48$) showed a 5.0% improvement in acceptance over the base model ($M=4.17$, $SD=1.46$), validating the practical applicability of our framework’s outputs.
Consistency. Significant differences emerged across models ($F(4,32)=3.73$, $p=0.013$, $\eta^{2}=0.318$). Grok 4 ($M=5.32$, $SD=1.27$) and OpenAI GPT-5 ($M=5.29$, $SD=1.31$) achieved the highest consistency ratings. DesignerLLM ($M=4.39$, $SD=1.36$) exhibited 8.1% better consistency than the base model ($M=4.06$, $SD=1.64$), demonstrating more reliable message convey capability.

### v2

@@ 3.5.1 Importance-based Scenario Sampling
For each scenario $i$, the shared action database stores tuples $(i,m,K_{i}^{m})$, where $m$ indexes the source model/round and $K_{i}^{m}$ is the kernel score. From these records, we compute $K_{i}^{\text{best}}=\max_{m}K_{i}^{m}$, $K_{i}^{\text{worst}}=\min_{m}K_{i}^{m}$, $\Delta K_{i}=K_{i}^{\text{best}}-K_{i}^{\text{worst}}$, and the candidate count $N_{i}$. We define the importance score as follows:
@@
  • Low best scores $K_{i}^{\text{best}}$ (room for improvement);
  • Small gaps $\Delta K_{i}$ (limited diversity);
  • High worst scores $K_{i}^{\text{worst}}$ (easier to improve);
  • Fewer previous sampling attempts.
@@ 3.5.2 Diverse Action Generation
For each sampled scenario, we generate diverse eHMI action sequences with the current fine-tuned model using diverse beam search Vijayakumar et al. (2016) (six beam groups with one beam each). Each output is validated against modality-specific JSON schema (eyes/arm/lightbar) via Pydantic, and invalid formats are discarded. We then render and evaluate the valid actions using the same pipeline as the initial database, compute their kernel scores, and append them to the shared action database. Repeating this step iteratively expands the candidate set and improves both diversity and quality over training rounds.
@@ 3.5.3 Preference Pair Construction
After expanding the shared score database, we form DPO preference pairs $(y^{+},y^{-})$ per scenario, where $y^{+}$ is a higher-quality action than $y^{-}$. This setup is analogous to the ordering of human preferences, as discussed in previous work Xia et al. (2025). Pair extraction is two-stage. (1) Max–min: for each scenario $i$, select the highest and lowest-scoring actions,
@@ 4.1 Experimental Settings
All fine-tuning processes use ShareGPT-formatted data Chiang et al. (2023) with an 80/20 train/test split. We fine-tune with LoRA Hu et al. (2022) in bfloat16: 3 epochs for Format-aware Fine-tuning and 1 epoch for DPO (to reduce overfitting, similar to previous works Bai et al. (2022b); Touvron et al. (2023); Liu et al. (2024)). DPO uses weight=1, sigmoid loss, and a preference coefficient of 0.01.
@@
VLM rating achieves approximately 90 evaluations per second with 8 threads. Single-thread inference requires about 7 seconds per scenario for the 7B model and 6 seconds per scenario for the 1.5B model. The rendering takes roughly 45 seconds per clip, operating in batch parallelization. Fine-tuning duration is approximately 3 hours for the 7B model and 1 hour 10 minutes for the 1.5B model in Format-aware Fine-tuning, and around 5 hours (7B) or 3 hours 30 minutes (1.5B) per DPO round. Excluding dataset construction and assuming three DPO rounds, the total co-learning time amounts to at least 60 hours for the 7B model and 38 hours for the 1.5B model. Importantly, this rendering-and-evaluation cost is incurred only during offline refinement. At deployment time, See2Refine uses the trained DesignerLLM to generate actions directly, without requiring video rendering or VLM-based evaluation.
@@ 4.2.3 Metric-wise Analysis
The metrics choices are inherited from previous eHMI work, but their effectiveness for VLM raters is unclear. Therefore, we analyze the VLM score distributions for three milestones: (1) the initial action database, (2) DesignerLLM-Apprentice-7B, and (3) DesignerLLM-7B.
As illustrated in Figure 7 in the Appendix, metrics behave differently over training. Phase 2 metrics (User Acceptance, Consistency) improve most clearly: their means increase, and distributions tighten, suggesting that they are the most sensitive and reliable signals for VLM-based evaluation.
In contrast, Phase 1 metrics show weaker trends. Targeting improves slightly, while trust changes little and provides poor preference separation in the initial database, making it less useful for constructing training pairs. We suspect that this reflects two factors. First, trust is a higher-level subjective judgment than acceptance or consistency and may therefore be harder for current VLMs to approximate reliably. Second, our controlled renderings omit environmental cues, such as background context, lighting, and occlusion, which may be particularly important for trust-related judgments. This suggests that See2Refine is currently better suited to optimize more directly perceivable communication qualities than nuanced subjective impressions such as trust.
@@ 4.3 Human Preference Alignment
We conduct a user study to (1) test whether improvements transfer from VLM raters to people, and (2) examine potential gaps between VLM and human preferences. Details are provided in Appendix A.
Setup. We evaluate two common modalities (eyes, lightbar) on eight scenario-message pairs from prior work Xia et al. (2025). We compare five models: DesignerLLM-base, DesignerLLM, and three commercial LLMs (OpenAI GPT-5, Claude Opus 4.1, Grok 4). Each model generated actions that are rendered as 1080p videos at 12 FPS with full scene context (80 videos total).
Participants and measures. We recruit 18 participants (21–30 years, M=26M{=}26, SD=2.7SD{=}2.7), split evenly by modality. For each scenario, participants rate videos in two stages: without the intended message (targeting, trust, perceived safety, mental workload) and with the intended message revealed (acceptance, consistency).
Results. In both modalities, DesignerLLM consistently outperform the base model and rank 4th overall (Figure 4). Relative to the base model, DesignerLLM improves the average scores by 7.7% (eyes) and 7.9% (lightbar), with the largest gains in targeting (15.4% eyes; 12.6% lightbar) and acceptance (5.0% eyes; 12.5% lightbar). Although participants show noticeable individual variation in how strictly they use the rating scale, the aggregated results exhibit a consistent relative advantage of DesignerLLM over the base model across most metrics. Commercial models achieve higher absolute ratings, but the improvements in our framework are statistically significant for most metrics, indicating that VLM-aligned training also yields measurable benefits for the quality of human-perceived eHMI.
@@ 5 Conclusion
We introduce See2Refine, leveraging VLM perceptual feedback as an automated approach to progressively enhance the capabilities of LLM-based eHMI action designers. We perform extensive experiments and explore methods to effectively expand the action dataset while preserving alignment with VLM preference performance. Additionally, we conduct a user study that shows that insights from VLM raters can also improve human perception scores. Finally, our work serves as an important practice in improving the message delivery capabilities of LLM-actuated systems. In future work, we plan to incorporate real human feedback and richer multimodal teaching signals to better capture subtle and diverse human preferences that are difficult to approximate with VLM raters alone.
@@
The current setup demonstrates that, with the See2Refine system, LLM-based eHMI action designers can achieve higher human perception scores, despite using VLM as an efficient substitute for real human annotators. However, as noted in previous work Xia et al. (2025), the preference order between a human and a VLM is consistent only when the difference in the rating score between two actions is sufficiently large. Replicating the nuanced preferences of individuals from different cultural backgrounds or age groups remains a challenge. Therefore, incorporating a real human annotator into the refinement process would be valuable.
There are two aspects worth exploring. First, effective fine-tuning of the VLM rater models could further improve system performance. However, collecting enough data for a single round of fine-tuning is still costly, so the potential marginal gains from such fine-tuning merit careful consideration. Second, rating alone is often not an information-rich method for annotation. If we consider our system as a simulation of a real teacher, employing a teaching approach, using rich cues such as voice, gestures, posture, and other non-verbal signals, could serve as a more natural and effective way to teach LLM designers accurate scenario-action pairs. This approach could also help reduce the ambiguity associated with preference pairs, enabling more effective distillation of human preferences into the LLM training process.
@@
In the current implementation (Section 3.1), we utilize generated scenarios and rendered action videos as visual cues for VLM raters. This approach has proven to be both effective and efficient. However, it focuses mainly on simulating the main elements of typical traffic scenarios, such as automated vehicles, while overlooking other factors that could influence visual perception outputs for both VLM raters and human road users, such as lighting conditions, obstructions, and other environmental factors. Therefore, the challenge is to incorporate more visual elements that accurately simulate real-world scenarios in rendered videos while maintaining reasonable rendering times to enhance the robustness of LLM-based action designers in practical applications. One potential trade-off solution is to combine different levels of visual detail within the same training session. For example, including videos with rich visual elements for only a small portion of the training data could improve robustness while keeping overall training costs relatively low.
Acknowledgments
This work was supported by a Canon Research Fellowship and JST CRONOS, Grant Number JPMJCS24K8, Japan
@@ A.1 Study Design and Procedure
We selected two eHMI modalities (lightbar and eyes) as simple and common modalities for the user study. Eight traffic scenario–message pairs were adopted from prior work Xia et al. (2025) covering common interactions between AVs and other road users. Five LLMs were tested: two 7B-parameter models (DesignerLLM-base, trained after Format-aware Fine-tuning, and DesignerLLM, after Iterative Preference-based Co-learning) and three state-of-the-art models (OpenAI GPT-5, Claude Opus 4.1, and Grok 4). Each model generated eHMI actions that were rendered as 1080p (1980×1080) videos at 12 FPS to ensure smooth playback for participants in Blender. Unlike the videos provided to the VLM raters (Section 3.2.2), these videos include rich contextual information. A total of 80 videos were created. Rendering all scenarios with detailed meshes and visual effects takes approximately 10 hours of computation.
@@
The user study was implemented and deployed using Gradio. Separate interfaces were created for the two eHMI modalities. Each interface consisted of five components: (1) a welcome page, (2) an introduction to general concepts of eHMI and the specific modality, including two demonstration videos, (3) a demographic information section, (4) main rating pages for eight pairs of scenarios and messages, and (5) an ending page. The interface included a session-based resume function, allowing participants to continue later using their assigned ID.
@@
In the demographic section, participants anonymously reported their age, gender, and familiarity with eHMIs. For the main user rating pages, eight scenarios are evaluated across different pages, each presented in two stages: the first stage, where the intended messages are not provided, and the second stage, where the intended messages are included. This results in a total of 16 pages. On each page, five videos are shown in randomized order. As illustrated in Figure 5, each video assessment follows a horizontal layout, with the video on the left and the questionnaires on the right. To ensure that no scores are missed, a verification mechanism is in place: participants must complete the current page before proceeding to the next, ensuring the completeness of the questionnaires. Each session took approximately 5 minutes to complete.
@@ A.2 Participants
We recruited 18 participants aged 21 to 30 years (mean age: 26, SD = 2.7), consisting of nine males and nine females. They were evenly assigned to evaluate the lightbar and eye modalities. Among the participants, one was very familiar with eHMIs, having extensive knowledge or experience; two had some familiarity, having heard of eHMIs before; and the remaining fifteen were unfamiliar with eHMIs. All participants received a \$10 gift card as compensation.
@@ A.3 Measurements
Without Intended Messages. Participants rated their initial impressions of the eHMI using four metrics:
  • Message Targeting (1–9): The extent to which the eHMI message was directed towards the participant.
  • Trust (1–9): The participant’s confidence in and willingness to rely on eHMI communication.
@@
  • Mental Workload (1–20): The cognitive effort required to interpret eHMI actions.
With Intended Messages. After revealing the intended messages, participants reassessed using two metrics:
  • User Acceptance (1–9): The willingness to accept an eHMI in daily life.
  • Consistency (1–9): The perceived alignment between the eHMI motion displayed and its intended message.
@@ A.4 Result Analysis
Prior to analysis, normality was assessed for these participant-level means using the Shapiro–Wilk test, and no serious deviations were found ($p>0.05$) for either modalities. Because the study used a within-subjects design with five model conditions, each metric was analyzed using repeated-measures ANOVA.
@@ A.4.1 Reliability Analysis
We assessed the reliability of subjective ratings before analyzing the effects of the model. Cronbach’s alpha was used to evaluate internal consistency within each metric across the five models. Internal consistency was excellent for both modalities (eye: $\alpha=0.925$; lightbar: $\alpha=0.953$), with all metrics exceeding $\alpha=0.90$. Inter-rater reliability, estimated with the intraclass correlation coefficient (ICC), was low (eye: $ICC=0.131$; lightbar: $ICC=0.087$), indicating limited absolute agreement across participants. Because our analyses focus on relative comparisons within each subject in a single session, high internal consistency is the primary requirement; the low ICC limits generalization to settings that require stable scores between raters or sessions Koo and Li (2016).
@@ A.4.2 Eye Modality Performance
For the eye modality condition with 9 participants, repeated measures ANOVA revealed significant main effects of the model type across all six metrics:
Message Targeting. ANOVA showed a significant main effect of model type ($F(4,32)=5.41$, $p=0.002$, $\eta^{2}=0.403$). Grok 4 ($M=5.47$, $SD=0.89$) and OpenAI GPT-5 ($M=5.47$, $SD=0.88$) achieved the highest ratings, while the base model Qwen2.5-7B-Instruct showed the lowest performance ($M=4.21$, $SD=1.65$). Notably, DesignerLLM ($M=4.86$, $SD=1.21$) demonstrated a 15.4% improvement over the base model, indicating that our co-learning framework successfully enhanced the message targeting metric despite its modest 7B parameter size.
Trust. A significant main effect was observed ($F(4,32)=4.66$, $p=0.005$, $\eta^{2}=0.368$). Grok 4 ($M=5.43$, $SD=1.04$) achieved the highest trust ratings, followed by OpenAI GPT-5 ($M=5.32$, $SD=1.15$). DesignerLLM ($M=4.65$, $SD=1.32$) achieved a 7.9% improvement over the base model ($M=4.31$, $SD=1.64$), demonstrating the framework’s effectiveness in generating more trustworthy eHMI action design.
@@
User Acceptance. ANOVA indicated a significant main effect ($F(4,32)=4.23$, $p=0.007$, $\eta^{2}=0.346$). Grok 4 ($M=5.39$, $SD=1.19$) and OpenAI GPT-5 ($M=5.31$, $SD=1.28$) achieved the highest acceptance ratings. DesignerLLM ($M=4.38$, $SD=1.48$) showed a 5.0% improvement in acceptance over the base model ($M=4.17$, $SD=1.46$), validating the practical applicability of our framework’s outputs.
Consistency. Significant differences emerged across models ($F(4,32)=3.73$, $p=0.013$, $\eta^{2}=0.318$). Grok 4 ($M=5.32$, $SD=1.27$) and OpenAI GPT-5 ($M=5.29$, $SD=1.31$) achieved the highest consistency ratings. DesignerLLM ($M=4.39$, $SD=1.36$) exhibited 8.1% better consistency than the base model ($M=4.06$, $SD=1.64$), demonstrating more reliable message-conduction capability.